OpenStudy (anonymous):

If g(x) = x2 + 2, find g(3). 9 8 11 6

OpenStudy (anonymous):

please help!!

OpenStudy (ipwnbunnies):

\[g(x) = x^2 + 2\] \[g(3) = (3)^2 + 2\]

OpenStudy (ipwnbunnies):

Just plug in 3 for x. :3

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

is the answer 11?

OpenStudy (ipwnbunnies):

Yas.

OpenStudy (anonymous):

thanks soo much